OpenStudy (anonymous):

Find the volume of a cone with a base radius of 3ft and a height of 10ft . Write the exact volume in terms of π , and be sure to include the correct unit in your answer.

OpenStudy (anonymous):

30π

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

\(\sf V = \dfrac{1}{3} \pi r^2 h\)

OpenStudy (igreen):

\(\sf 30 \pi\) is correct

OpenStudy (igreen):

@Cassandra_Lea_96

OpenStudy (anonymous):

awesome =0 so I just put ft at the end and i'm good?

OpenStudy (anonymous):

@iGreen